# SoliReward: Mitigating Susceptibility to Reward Hacking and Annotation Noise in Video Generation Reward Models

### v2

@@ 2.3 Post-Training on Visual Generation
Group Relative Policy Optimization. Adapted for visual generation, GRPO-based online algorithms [36, 16, 13, 19] addresses the instability of standard policy gradient methods. They utilize online feedback from reward model as guidance for optimization. By introducing Brownian motion, the ordinary differential equation (ODE) inherent in flow matching [15] is transformed into a stochastic differential equation (SDE). This infusion of stochasticity enhances the model’s exploration within the generation space.
@@
Reward Feedback Learning. ReFL [35, 4, 23] is an alignment technique that utilizes a differentiable reward model, employing the reward signal directly as the loss function. This mechanism allows for end-to-end optimization by backpropagating gradients from the reward model directly through the generative model’s parameters, thus avoiding the complexities of GRPO.
@@ 3 Methodology
In this section, we detail our pipeline. As show in Fig. 1, we first introduce our data annotation approach (Sec. 3.1) and cross-prompt pairing strategy (Sec. 3.2). We then describe the training loss, a win-tie variant of the Bradley-Terry model (Sec. 3.3) designed to mitigate reward hacking and the reward model architecture HPQA (Sec. 3.4).
@@ 3.1 Data Annotation
The efficacy of an RM depends heavily on preference data quality, yet standard pairwise and point-wise methods suffer from high annotator ambiguity and disagreement [9, 17]. To mitigate this, we design a single-item binary annotation. By evaluating single videos against objective criteria (Pass/Fail), we diminish label noise and improve IAA, as shown in Sec. 4.1.
We focus on three critical dimensions: Physical Plausibility, Subject Deformity, and Semantic Alignment. This binary scheme yields high-confidence signals. We annotate $250\mathrm{k}$ in-house training videos and a $50\mathrm{k}$-video out-of-distribution (OOD) test set generated by other SOTA models. These samples originate from $20\mathrm{k}$ unique prompts (please refer to Appendix Sec. 8 for details).
@@ 3.4 Hierarchical Progressive Query Attention
The operational flow begins by generating an initial query vector $q^{(1)}\in\mathbb{R}^{B\times 1\times D}$. This is achieves by attending to the first specified hidden state $H_{l_{1}}$ using a single, learnable query vector $q^{(0)}\in\mathbb{R}^{1\times 1\times D}$. This vector $q^{(0)}$ acts as the Query, while the input hidden state $H_{l_{1}}$ serves as both the Key and Value in a standard Multi-Head Attention (MHA) operation:
@@ 5 Conclusion and Future Work
Limitation & Future Work. We primarily validate the reward model’s effectiveness in the post-training phase of text-to-video (T2V) generation. Future work can be extended to other conditional generation tasks, such as image-to-video (I2V). Besides, our reward model scores input videos along a single dimension, and multiple dimensions can be fused into a single RM by designing several learnable queries. We observe more performance gains when scaling our RM from 1B to 8B parameters, but diminishing returns when scaling further to 14B. This finding on model scaling highlights a key avenue for future exploration.
@@ 7.3.1 Physical Plausibility and Subject Deformity
We select corresponding VBench2 Human Fidelity, including human anatomy, human clothes and human identity, as the evaluation dimensions. As detailed in  Tab. 10, our RM outperforms both the baseline and the VideoAlign-MQ guided training.
Table 10: Comparison of different approaches on human preference benchmarks on VBench2 [39]. The metrics are: Anatomy (Human Anatomy), Clothes (Human Clothes), Identity (Human Identity), and Fidelity (Human Fidelity).
Backbone RM Anatomy Clothes Identity Fidelity HunyuanVideo - 0.8915 0.8905 0.7457 0.8426 HunyuanVideo MQ 0.9009 0.8714 0.8361 0.8695 HunyuanVideo Ours 0.9164 0.9231 0.8601 0.8999
@@ 7.3.2 Semantic Alignment
Table 11: Implementation details for reward model training.
Table 12: Implementation details for post-training.
@@ 10.2 Post-Training
We conducted post-training experiments on a 5-node CentOS cluster, each node featuring 8 NVIDIA H800 GPUs (80GB CUDA memory) and an Intel Xeon Platinum 8476C Processor. Other key hyperparameters are detailed in Tab. 12.

### v3

@@ 2.3 Post-Training on Visual Generation
Group Relative Policy Optimization. Adapted for visual generation, GRPO-based online algorithms [38, 16, 13, 19, 42] addresses the instability of standard policy gradient methods. They utilize online feedback from reward model as guidance for optimization. By introducing Brownian motion, the ordinary differential equation (ODE) inherent in flow matching [15] is transformed into a stochastic differential equation (SDE). This infusion of stochasticity enhances the model’s exploration within the generation space.
@@
Reward Feedback Learning. ReFL [37, 4, 24, 21] is an alignment technique that utilizes a differentiable reward model, employing the reward signal directly as the loss function. This mechanism allows for end-to-end optimization by backpropagating gradients from the reward model directly through the generative model’s parameters, thus avoiding the complexities of GRPO.
@@ 3 Methodology
In this section, we detail our pipeline. As shown in Fig. 1, we first introduce our data annotation approach (Sec. 3.1) and cross-prompt pairing strategy (Sec. 3.2). We then describe the training loss, a win-tie variant of the Bradley-Terry model (Sec. 3.3) designed to mitigate reward hacking and the reward model architecture HPQA (Sec. 3.4).
@@ 3.1 Data Annotation
The efficacy of an RM depends heavily on preference data quality, yet standard pairwise and point-wise methods suffer from high annotator ambiguity and disagreement [9, 17]. To mitigate this, we design a single-item binary annotation. By evaluating single videos against objective criteria (Pass/Fail), we diminish label noise and improve IAA, as shown in Sec. 4.1. We focus on three critical dimensions: Physical Plausibility, Subject Deformity, and Semantic Alignment. This binary scheme yields high-confidence signals. We annotate $250\mathrm{k}$ in-house training videos and a $50\mathrm{k}$-video out-of-distribution (OOD) test set generated by other SOTA models. These samples originate from $20\mathrm{k}$ unique prompts (please refer to Appendix Sec. 8 for details).
@@ 3.4 Hierarchical Progressive Query Attention
The operational flow begins by generating an initial query vector $q^{(1)}\in\mathbb{R}^{B\times 1\times D}$. This is achieved by attending to the first specified hidden state $H_{l_{1}}$ using a single, learnable query vector $q^{(0)}\in\mathbb{R}^{1\times 1\times D}$. This vector $q^{(0)}$ acts as the Query, while the input hidden state $H_{l_{1}}$ serves as both the Key and Value in a standard Multi-Head Attention (MHA) operation:
@@ 5 Conclusion and Future Work
Limitation & Future Work. Future work can be extended to other conditional generation tasks, such as image-to-video (I2V). Besides, our reward model scores input videos along a single dimension, and multiple dimensions can be fused into a single RM by designing several learnable queries. We observe more performance gains when scaling our RM from 1B to 8B parameters, but diminishing returns when scaling further to 14B. This finding on model scaling highlights a key avenue for future exploration.
@@ 7.2 Model Scaling
7.3 BT, BTT and BT-WT
7.4 HPQA Layer Indices
Empirical results in Table 11 and Table 12 demonstrate that the HPQA architecture maintains its efficacy across diverse VLM backbones. Furthermore, our ablation study indicates that aggregating representations from 4 to 5 intermediate layers yields better performance.
Table 11: Performance of Qwen2-VL (2B) [31] across different layer indices.
Task Layer Indices ID OOD ACC Margin ACC Margin Phy & Deform ‘Yes’ Token 75.40 1.71 77.05 1.69 {14, 28} 75.79 1.71 77.24 1.85 {9, 18, 28} 75.20 1.64 77.59 1.59 {7, 14, 21, 28} 76.34 1.72 77.78 1.81 {6, 12, 18, 24, 28} 76.29 1.77 78.89 1.84 {5, 10, 14, 19, 23, 28} 74.11 1.67 75.73 1.56 TA ‘Yes’ Token 77.33 3.41 55.28 0.70 {7, 14, 21, 28} 78.62 3.55 56.19 0.56 {5, 10, 14, 19, 23, 28} 77.53 3.17 56.33 0.54
Table 12: Performance of Qwen2.5-VL (3B) and InternVL3 (14B) across different layer indices.
Task Model Layer Indices ID OOD ACC Margin ACC Margin Phy & Deform Qwen2.5-VL (3B) ‘Yes’ Token 75.45 1.61 40.76 -1.52 {9, 18, 27, 36} 73.68 1.61 76.38 1.50 {7, 14, 21, 28, 36} 73.51 1.47 76.61 1.39 {6, 12, 18, 24, 30, 36} 72.87 1.43 74.27 1.35 InternVL3 (14B) {16, 32, 48} 78.10 3.56 80.99 5.39 {12, 24, 36, 48} 78.50 3.52 80.31 5.17 {9, 18, 27, 36, 48} 78.20 3.71 81.16 5.21 TA Qwen2.5-VL (3B) ‘Yes’ Token 76.19 1.89 27.73 0.24 {9, 18, 27, 36} 76.54 2.84 59.06 0.51 {6, 12, 18, 24, 30, 36} 77.23 2.80 58.16 0.35 scores degenerate to discrete values.
@@ 7.5.1 Physical Plausibility and Subject Deformity
We select corresponding VBench2 Human Fidelity, including human anatomy, human clothes and human identity, as the evaluation dimensions. As detailed in  Tab. 13, our RM outperforms both the baseline and the VideoAlign-MQ guided training. Both BT-WT and HPQA generate positive returns, with the former yielding even higher returns.
Table 13: Comparison of different approaches on human preference benchmarks from VBench2 [41]. The evaluation metrics include Anatomy (Human Anatomy), Clothes (Human Clothes), Identity (Human Identity), and Fidelity (Human Fidelity).
Method Anatomy Clothes Identity Fidelity (Overall) Baseline 0.8915 0.8905 0.7457 0.8426 VideoAlign-MQ 0.9009 0.8714 0.8361 0.8695 ‘Yes’ Token w/ BT 0.9154 0.9067 0.7778 0.8666 ‘Yes’ Token w/ BT-WT 0.9312 0.9259 0.8013 0.8861 HPQA w/ BT 0.9153 0.9064 0.7861 0.8693 HPQA w/ BTT 0.9254 0.8922 0.7923 0.8700 HPQA w/ BT-WT 0.9164 0.9231 0.8601 0.8999
@@ 9.2 Post-Training
We conducted post-training experiments on a 5-node CentOS cluster, each node featuring 8 NVIDIA H800 GPUs (80GB CUDA memory) and an Intel Xeon Platinum 8476C Processor. Key hyperparameters are detailed in Tab. 17.
@@ 10 More Visualization Results
Table 16: Implementation details for reward model training.
Table 17: Implementation details for post-training.